OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 5 points) [9.01] The graph of which quadratic equation is shown below?

OpenStudy (anonymous):

then the answers y = x2 + 4x − 3 y = x2 − 4x + 3 y = −x2 + 4x − 3 y = −x2 − 4x + 3

OpenStudy (anonymous):

I thought it be B am i right??

OpenStudy (anonymous):

yes...you are correct.